GPR, Inc., has an inventory turnover of 27.06 times, a payables turnover of 15.09 times, and a receivables turnover of 8.56 times. What is the length of the company's cash cycle? Assume 365 days per year.
Answer #1

Calculation of cash cycle:

Days inventory outstanding= 365/inventory Turnover ratio= 365/27.06= 13.49 days

Days sale outstanding= 365/receivable turnover ratio= 365/8.56= 42.64 days

Days payable outstanding= 365/payable turnover= 365/15.09= 24.19

Cash cycle= days inventory outstanding+ days sale outstanding- days payable outstanding

Cash cycle= 13.49+42.64-24.19= 31.94 days

Cash cycle= 31.94 days